OpenStudy (anonymous):

If a scale of a map is 1 inch: 75 milesand it is 3 1/2 inches from St.Louis, Missouri to indianapolis, Indiana, how long would it take to travel from one city to the other at 50 miles per hour?

OpenStudy (anonymous):

the real distance is 3.5*75=262.5 miles time needed=262.5/50=5.25 hours

OpenStudy (anonymous):

ghass solution is correct..